OpenStudy (happykiddo):

please show how to write this expression as a single logarithm::::: Give me a second to post the expression

OpenStudy (happykiddo):

\[-2\log_{3} (1/x)+1/3\log_{3}\sqrt{x} \]

OpenStudy (anonymous):

\[\log_b(A) + \log_b(B) = \log_b(AB)\]\[\log_b(A) - \log_b(B) = \log_b \left( \frac{ A }{ B } \right)\]\[\log_b(A^B) = B\ \log_b(A)\]

OpenStudy (anonymous):

use the third one first, then the first one

OpenStudy (happykiddo):

i got \[\log_{3}x^(1/6)\]

OpenStudy (ranga):

-2log(1/x) = log(1/x)^(-2) = log(1/x^(-2)) = log(x^2) 1/3log(sqrt(x)) = 1/3log(x)^(1/2) = log((x)^(1/2))^(1/3) = log(x)^(1/6) log(x^2) + log(x)^(1/6) = log(x^2 * x^(1/6)) = log(x)^(2+1/6) = log(x)^(13/6) = 13/6 * log(x) The log is to the base 3.

OpenStudy (happykiddo):

thank you : )

OpenStudy (ranga):

you are welcome :)